OpenStudy (anonymous):

evaluate the following logarithm: log3 9

myininaya (myininaya):

log(base3)9=(ln9)/(ln3)=[ln3^2]/ln3=2(ln3)/(ln3)=2 also think if log(base3)9=y then 3^y=9 so y=2

OpenStudy (anonymous):

thankyou